Electrical field

Three positive point charges of 3.0 nC, 6.0 nC, and 2.0 nC,respectively, are arranged in a triangular pattern, as shown. Findthe magnitude and direction of theelectrical force acting on the6.0 nC charge.
Answer #1
The three positive point charges are q1= 3.0nC,q2= 6.0 nC and q3= 2.0 nCThe electrical force acting on the charge q2 due tothe charge q1 isF1= (1/4πεo) *(q1q2/r2)Here,(1/4πεo) = 9 * 109Nm2/C2 and r = (1 + 1)1/2 =√2or F1=9 * 109* (3.0 *10-9* 6.0 *10-9/(√2)2)or F1= 81 * 10-9 NThe electrical force acting on the charge q2 due tothe charge q3 isF2= 9 * 109* (6.0 *10-9* 2.0 *10-9/(√2)2)or F2=54 * 10-9 NThe magnitude of the net electrical force acting on the 6.0 nCcharge isF = F1 + F2or F = 81 * 10-9 +54 * 10-9Nor F =135 * 10-9 NThe direction of the net electrical force acting on the 6.0 nCcharge istanα = (F1 * cosθ/F1 +F2 * sinθ)Here,sinθ = (1/√2)or θ = 45oor tanα = (81 * 10-9 * cos(45o)/81* 10-9 +54 * 10-9 *sin(45o))or tanα = 0.4807or α = 25.67oTherefore,the direction of the net electrical force acting onthe 6.0 nC charge is α = 25.67o counterclockwisefrom the positive x-axis.
answered by: <3